1 [1 07 0 2 2 0 0 3 Find the eigenvalues and eigenvectors of A=
Answer #1

Given A= 2 2. 3 1-2 - lara Il= 2-2 2 3-2 =(1-2) (2-2) (3-2)=0 => 2=1,2,3 are eigenvalues of Now when 2=1 xi CA-I) X=0 where